A parking -meter slot receives dimes and nickels. When emptied, the box produced 70 coins worth $4.85. How many nickels and how many dimes were there?
Answer by ptaylor(1894)


WE CAN WORK THIS PROBLEM EITHER USING ONE EQUATION AND ONE UNKNOWN OR TWO EQUATIONS AND TWO UNKNOWNS. WE'LL SOLVE IT BOTH WAYS:
One unknown:
Let x=number of nickels
Then 70-x=number of dimes
Let's deal in pennies
5x=value of nickels
10(70-x)=700-10x=value of dimes
We are told that the above amounts add up to 485, so:
5x+700-10x=485 subtract 700 from each side
5x-10x+700-700=485-700 collect like terms
-5x=-215 divide each side by -5
x=43-------------------------------number of nickels
70-x=70-43=27--------------------------number of dimes
CK
43+27=70
70=70
and
5*43+10*27=485
215+270=485
485=485
Two unknowns:
Let x=number of nickels
And y=number of dimes
x+y=70------------------------------eq1
5x+10y=485--------------------------eq2
From eq1, y=70-x. Substitute this into eq2
5x+10(70-x)=485-----------------------same equation as before

Does this help???-------------ptaylor